OpenStudy (anonymous):

Convert 0.625g/L to g/cm

OpenStudy (mathstudent55):

Do you mean Convert \(~0.625 \dfrac{g}{L}~\) to \(~\dfrac{g}{cm^3} \) ?

hero (hero):

\[0.625 \frac{g}{L} = x \frac{g}{cm^3}\]\[\frac{0.625}{L} = \frac{x}{cm^3}\]\[\frac{0.625}{1000 cm^3} = \frac{x}{cm^3}\]\[\frac{0.625}{1000} = x\]

OpenStudy (mathstudent55):

You need to know that \(1 L = 1000 ~mL\) and \(1~mL = 1 ~cm^3 \therefore 1~ L = 1000 ~cm^3\) \(0.625 \dfrac{g}{L} \times \dfrac{1~L}{1000~cm^3} = \dfrac{0.625}{1000} \dfrac{g}{cm^3} \)

OpenStudy (mathstudent55):

Not correct. This is correct: \(1~L = 1000~mL = 1000~cm^3\)

hero (hero):

That's what I meant.

OpenStudy (mathstudent55):

Ok, I figured, since you did the problem well.

hero (hero):

I did it my way.